# Achieving Linear Speedup with ProxSkip in Distributed Stochastic Optimization

### v3

@@ Abstract
The ProxSkip algorithm for distributed optimization is gaining increasing attention due to its proven benefits in accelerating communication complexity while maintaining robustness against data heterogeneity. However, existing analyses of ProxSkip are limited to the strongly convex setting and do not achieve linear speedup, where convergence performance increases linearly with respect to the number of nodes. So far, questions remain open about how ProxSkip behaves in the non-convex setting and whether linear speedup is achievable. In this paper, we revisit distributed ProxSkip and address both questions. We demonstrate that the leading communication complexity of ProxSkip is $\mathcal{O}(\nicefrac{{p\sigma^{2}}}{{n\epsilon^{2}}})$ for non-convex and convex settings, and $\mathcal{O}(\nicefrac{{p\sigma^{2}}}{{n\epsilon}})$ for the strongly convex setting, where $n$ represents the number of nodes, $p$ denotes the probability of communication, $\sigma^{2}$ signifies the level of stochastic noise, and $\epsilon$ denotes the desired accuracy level. This result illustrates that ProxSkip achieves linear speedup and can asymptotically reduce communication overhead proportional to the probability of communication. Additionally, for the strongly convex setting, we further prove that ProxSkip can achieve linear speedup with network-independent stepsizes.
Index Terms:
Distributed optimization, stochastic non-convex optimization, local update, linear speedup.
@@ I Introduction
In this work, we consider the following distributed stochastic optimization problem by a group of agents $[n]:=\{1,2,\ldots,n\}$ connected over a network:
$f^{\star}=\min_{{\bf{x}}\in\mathbb{R}^{d}}\Big{[}f({\bf{x}}):=\frac{1}{n}\sum_{i=1}^{n}f_{i}({\bf{x}})\Big{]},$ (1)
with $\text{with }f_{i}({\bf{x}})=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F_{i}({\bf{x}},\xi_{i})],$
@@
TABLE I: Comparisons with existing convergence rates of ProxSkip for distributed optimization. $\sigma^{2}$ is the variance of the stochastic gradient. $\zeta_{0}:=\max\{1-\alpha\mu,1-p^{2}\}$, the definition of $\zeta_{{\mathrm{new}}}$ and $\zeta$ can be found in [48] and Theorem 1, respectively. CVX, N-CVX, and S-CVX mean convex, non-convex, and strongly convex, respectively.
@@
In this paper, we revisit ProxSkip for distributed stochastic optimization and provide answers to both questions. Specifically, we develop a new analysis with a novel proof technique under non-convex, convex, and strongly convex settings. Through this analysis, we obtain several new results that are comparable to the bounds of state-of-the-art distributed algorithms while achieving linear speedup bounds. We highlight our contributions as follows:
@@
    (N-)CVX: $\mathcal{O}\bigg{(}\frac{1}{\alpha T}+\frac{\alpha^{2}}{(1-\lambda_{2})^{2}T}+\frac{\alpha\sigma^{2}}{n}+\frac{\sigma^{2}\alpha^{2}}{1-\lambda_{2}}\bigg{)},$
    S-CVX: $\mathcal{O}\bigg{(}(1-\alpha\mu)^{T}a_{0}+\frac{\alpha\sigma^{2}}{\mu n}+\frac{\sigma^{2}\alpha^{2}}{\mu(1-\lambda_{2})}\bigg{)},$
    where $\alpha$ is the stepsize of ProxSkip, $\sigma^{2}$ denotes the variance of the stochastic gradient, $1-\lambda_{2}$ is a topology-dependent quantity that approaches 00 for a large and sparse network, $\mu$ is the strongly convex constant, and $a_{0}$ is a constant that depends on the initialization. To the best of our knowledge, it is the first work that establishes the convergence rate of probabilistic distributed methods for non-convex settings. We offer a comparison of convergence rates of ProxSkip for (1) in Table I.
@@
  • We elucidate the effects of noise, local steps, and data heterogeneity on the convergence of ProxSkip in stochastic non-convex, convex, and strongly convex settings. We demonstrate the robustness of ProxSkip against data heterogeneity while enhancing communication efficiency by local updates. Furthermore, we show that the convergence rates exhibited by ProxSkip in stochastic settings are comparable with those of existing state-of-the-art distributed algorithms incorporating local updates [31, 35, 36, 37, 42] (see Table II).
@@ II-A Network graph
The mixing matrix $\mathbf{W}\in[0,1]^{n\times n}$ is symmetric, doubly stochastic, and primitive. Let $\lambda_{1}=1$ denote the largest eigenvalue of the mixing matrix $\mathbf{W}$, and the remaining eigenvalues are denoted as $11>\lambda_{2}\geq\lambda_{3}\geq\cdots\geq\lambda_{n}>-$ 1.
@@ II-B Algorithm description
Here, $\alpha>0$ is the stepsize, $\beta>0$, $\mathbf{G}^{t}=[\mathbf{g}_{1}^{t},\mathbf{g}_{2}^{t},\ldots,\mathbf{g}_{n}^{t}]^{\sf T}\in\mathbb{R}^{n\times d}$ with $\mathbf{g}_{i}^{t}$ representing the stochastic gradient of $\nabla f_{i}(\mathbf{x}_{i}^{t})$, $\theta_{t}=1$ with probability $p$ and $\theta_{t}=0$ with probability $1-p$, and $\mathbf{Y}^{t}$ is the control variate. At each iteration $t\geq 0$, communication takes place with a probability $p\in(0,1]$. In the absence of communication, the update $\mathbf{X}^{t+1}=\mathbf{X}^{t}-\alpha\mathbf{G}^{t}-\mathbf{Y}^{t}$ is performed, while $\mathbf{Y}^{t+1}$ remains unchanged. This allows for multiple iterations of local computations to be performed between communication rounds. Decomposing the updates for individual nodes, we provide a detailed implementation in Algorithm 1.
@@
2:  Flip coins $[\theta_{0},\ldots,\theta_{T-1}]$, where $\theta_{t}\in\{0,1\}$, with $\mathop{{\bf{P}}}(\theta_{t}=1)=p$.
@@
5:     $\hat{{\bf{z}}}^{t}_{i}={\bf{x}}^{t}_{i}-\alpha{\bf{g}}^{t}_{i}-{\bf{y}}^{t}_{i}$triangleright update the prediction variate $\hat{{\bf{z}}}^{t}_{i}$
@@
7:        ${\bf{x}}^{t+1}_{i}=\sum_{j=1}^{n}\widetilde{W}_{ij}\hat{{\bf{z}}}^{t}_{j}$triangleright communicate with probability $p$
8:        ${\bf{y}}^{t+1}_{i}={\bf{y}}^{t}_{i}+\beta(\hat{{\bf{z}}}^{t}_{i}-{\bf{x}}^{t+1}_{i})$triangleright update the control variate ${\bf{y}}^{t+1}_{i}$
@@
10:        ${\bf{y}}^{t+1}_{i}={\bf{y}}^{t}_{i},~{}{\bf{x}}^{t+1}_{i}=\hat{{\bf{z}}}^{t}_{i}$triangleright skip communication
@@ III-A Preliminary
When $\sigma^{2}=0$, by setting $\alpha=\nicefrac{{1}}{{L}}$ and $\chi=1$, we can deduce from (3) that the communication complexity of ProxSkip to achieve $\epsilon$-accuracy, i.e., $\mathbb{E}[|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}]\leq\epsilon$, is given by $\mathcal{O}((p\kappa+\nicefrac{{1}}{{p(1-\lambda_{2})}})\mathrm{log}\nicefrac{{1}}{{\epsilon}})$, where $\kappa=\nicefrac{{L}}{{\mu}}$. If the network is sufficiently well-connected, i.e., $\nicefrac{{1}}{{(1-\lambda_{2})\kappa}}<1$, and we set $p=\sqrt{\nicefrac{{1}}{{(1-\lambda_{2})\kappa}}}$, the iteration complexity becomes $\mathcal{O}(\sqrt{\nicefrac{{\kappa}}{{1-\lambda_{2}}}}\ \mathrm{log}\nicefrac{{1}}{{\epsilon}})$, achieving the optimal communication complexity as proven by [50].
When $\sigma^{2}\neq 0$, based on (3) and the fact that $\frac{\alpha^{2}\sigma^{2}}{1-\zeta}=\frac{\alpha^{2}\sigma^{2}}{\alpha\mu}=\mathcal{O}(\alpha\sigma^{2})$, we can conclude that the local solution ${\bf{x}}_{i}^{t}$ generated by ProxSkip converges to the global minimizer ${\bf{x}}^{\star}$ at a linear rate until it reaches an $\mathcal{O}(\alpha\sigma^{2})$-neighborhood of ${\bf{x}}^{\star}$. However, it is important to note that relying solely on equation (3) is not sufficient to achieve the desired linear speedup term $\mathcal{O}(\frac{\alpha\sigma^{2}}{n})+\mathcal{O}(\alpha^{2})$. This indicates that the direct extension of the analysis techniques proposed in [43] or [49] to the stochastic scenario does not guarantee linear speedup, despite ensuring convergence. Therefore, further analysis is required to achieve the desired linear speedup.
@@ III-B Main theorem—Convergence rate of ProxSkip
Suppose that Assumptions 1, 2, and 4 hold. Let $\bar{{\bf{x}}}^{t}=\frac{1}{n}\sum_{i=1}^{n}{\bf{x}}_{i}^{t}$ denote the iterates of Algorithm 1 and ${\bf{x}}^{\star}$ solves (1). For sufficiently small $\alpha$, $\chi=\mathcal{O}(\max\{1,\nicefrac{{(1-p)}}{{(1-\lambda_{2})}}\})$, and $\beta=1$, we have the following convergence results.
@@
$\leq\mathcal{O}\bigg{(}\underbrace{\frac{F_{0}}{\alpha T}+\frac{\chi^{2}L^{2}\varsigma^{2}_{0}\alpha^{2}}{(1-\lambda_{2})^{2}T}}_{{\mathrm{deterministic\ part}}}+\underbrace{\frac{\alpha\sigma^{2}L}{n}+\frac{\chi L^{2}\sigma^{2}\alpha^{2}}{1-\lambda_{2}}}_{{\mathrm{stochastic\ part}}}\bigg{)}.$ (4)
@@
$\leq\mathcal{O}\bigg{(}\underbrace{\frac{R_{0}^{2}}{\alpha T}+\frac{\chi^{2}L\varsigma^{2}_{0}\alpha^{2}}{(1-\lambda_{2})^{2}T}}_{{\mathrm{deterministic\ part}}}+\underbrace{\frac{\alpha\sigma^{2}}{n}+\frac{\chi L\sigma^{2}\alpha^{2}}{1-\lambda_{2}}}_{{\mathrm{stochastic\ part}}}\bigg{)}.$ (5)
@@
$\mathbb{E}\!\left[\big{\|}\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\big{\|}^{2}\right]\leq$ $\underbrace{\Big{(}1-\frac{\alpha\mu}{4}\Big{)}^{T}\Big{(}\big{\|}\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big{\|}^{2}+\frac{8\chi\alpha^{2}\varsigma_{0}^{2}}{1-\lambda_{2}}\Big{)}}_{{\mathrm{deterministic\ part}}}$
$+\mathcal{O}\bigg{(}\underbrace{\frac{\alpha\sigma^{2}}{\mu n}+\frac{\chi L\sigma^{2}\alpha^{2}}{\mu(1-\lambda_{2})}}_{{\mathrm{stochastic\ part}}}\bigg{)}.$ (6)
For the non-convex setting, Theorem 1 demonstrates that the ProxSkip algorithm converges to a radius around some stationary point. Without any additional assumptions, a stationary point is the best guarantee possible and is a satisfactory criterion to measure the performance of distributed methods with nonconvex objectives [31]. For the convex case, Theorem 1 shows that ProxSkip converges around some optimal solution. When $\sigma^{2}=0$, i.e., in the deterministic case, ProxSkip converges exactly with sublinear and linear rates for N-CVX/CVX and S-CVX settings, respectively.
Note that stochastic part in convergence rates (1), (1), and (1), which all can be rewritten as $\mathcal{O}(\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2})$. It follows from Theorem 1 that, in non-convex, convex, and strongly convex settings, the convergence rate of ProxSkip is
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}\right]\leq\mathcal{O}\left(\frac{1}{\alpha T}\right)+\mathcal{O}\left(\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right),$
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[f(\bar{{\bf{x}}}^{t})-f^{\star}\right]\leq\mathcal{O}\left(\frac{1}{\alpha T}\right)+\mathcal{O}\left(\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right),$
$\mathbb{E}\!\left[\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\|^{2}\right]\leq\mathcal{O}\left((1-\alpha\mu)^{T}\right)+\mathcal{O}\left(\frac{\alpha\sigma^{2}}{n}+\alpha^{2}\sigma^{2}\right),$
respectively. Thus, it is established in Theorem 1 that the linear speedup term $\mathcal{O}(\frac{\alpha\sigma^{2}}{n})$ can be achieved. When the stepsize is sufficiently small the term $\mathcal{O}(\frac{\alpha\sigma^{2}}{n})$ dominates convergence rates (1), (1), and (1), which improve linearly with the number of nodes $n$.
Setting $\alpha=\sqrt{\frac{n}{T}}$ for sufficiently large $T$ for non-convex and convex settings, it holds that the rates are bounded by
$\mathcal{O}\left(\sqrt{\frac{1+\sigma^{2}}{nT}}+\frac{n\chi\sigma^{2}}{(1-\lambda_{2})T}+\frac{n\chi^{2}}{(1-\lambda_{2})^{2}T^{2}}\right)=\mathcal{O}\left(\sqrt{\frac{1}{nT}}\right).$
For the strongly convex setting, letting $\alpha=\frac{4\ln T^{2}}{\mu T}$ for sufficiently large $T$, it holds that $1-\frac{\alpha\mu}{4}\leq{\mathrm{exp}}(-\frac{\alpha\mu T}{4})=\frac{1}{T^{2}}$, where ${\mathrm{exp}}(\cdot)$ denotes the exponential function, thus the rate is bounded by
$\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{nT}+\frac{1}{T^{2}}+\frac{\chi\sigma^{2}}{(1-\lambda_{2})T^{2}}+\frac{\chi}{(1-\lambda_{2})T^{4}}\right)=\tilde{\mathcal{O}}\left(\frac{1}{nT}\right).$
When $T$ is sufficiently large, the term $\frac{1}{\sqrt{nT}}$ (or $\frac{1}{nT}$ for the strongly convex setting) will dominate the rate. In this scenario, ProxSkip requires $T=\Omega\left(\frac{1}{n\epsilon^{2}}\right)$ (or $T=\Omega\left(\frac{1}{n\epsilon}\right)$) iterations to reach a desired $\epsilon$-accurate solution, thus the convergence accuracy improves linearly with $n$.
In addition, based on Theorem 1, we can even get a tighter rate by carefully selecting the stepsize to obtain the following result.
@@
$\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p\sqrt{\chi}}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{\chi}{(1-\lambda_{2})\epsilon}\right)$ (7)
@@
$\mathcal{O}\left(\frac{p\sigma^{2}}{n\epsilon^{2}}+\frac{p\sqrt{\chi}}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{\chi}{(1-\lambda_{2})\epsilon}\right)$ (8)
@@
$\tilde{\mathcal{O}}\left(\frac{p\sigma^{2}}{n\epsilon}+\frac{p\sqrt{\chi}}{\sqrt{1-\lambda_{2}}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\chi\mathrm{log}\nicefrac{{1}}{{\epsilon}}}{1-\lambda_{2}}\right)$ (9)
expected communication rounds. Here, the notation $\tilde{\mathcal{O}}(\cdot)$ ignores logarithmic factors.
@@
[…]

### v4

@@ Abstract
The ProxSkip algorithm for distributed optimization is gaining increasing attention due to its effectiveness in reducing communication. However, existing analyses of ProxSkip are limited to the strongly convex setting and fail to achieve linear speedup with respect to the number of nodes. Key questions regarding its behavior in the non-convex setting and the achievability of linear speedup remain open. In this paper, we revisit ProxSkip and address both questions. We provide a comprehensive analysis for stochastic non-convex, convex, and strongly convex problems, revealing the effects of gradient noise, local updates, network connectivity, and data heterogeneity on its convergence. We prove that ProxSkip achieves linear speedup across all three settings, and can further achieve linear speedup with network-independent stepsizes in the strongly convex setting. Moreover, we show that properly increasing local updates effectively reduces communication complexity.
@@ I Introduction
In this work, we consider the following distributed stochastic optimization problem by a group of agents $n]:=\{1,2,\ldots,n\}$ connected over a network:
$f^{\star}=\min_{{\bf{x}}\in\mathbb{R}^{d}}\Big[f({\bf{x}})=\frac{1}{n}\sum_{i=1}^{n}f_{i}({\bf{x}})\Big],$ (1)
$\text{with }f_{i}({\bf{x}})=\mathbb{E}_{\xi_{i}\sim\mathcal{D}_{i}}[F_{i}({\bf{x}},\xi_{i})],$
@@
TABLE I: Comparisons with existing convergence rates of ProxSkip for distributed optimization. $\sigma^{2}$ is the variance of the stochastic gradient. $\zeta_{0}:=\max\{1-\alpha\mu,1-p^{2}\}$, the definition of $\zeta_{{\mathrm{new}}}$ and $\zeta$ can be found in [48] and Theorem 1, respectively.
@@
In this paper, we revisit ProxSkip for distributed stochastic optimization and provide answers to both questions. Specifically, we develop a new analysis with a novel proof technique under non-convex (N-C), convex (C), and strongly convex (SC) settings. Through this analysis, we obtain several new results that are comparable to the bounds of state-of-the-art distributed algorithms while achieving linear speedup bounds. We highlight our contributions as follows:
@@
    (N-)C: $\mathcal{O}\left(\frac{\sigma}{\sqrt{nT}}+\frac{\chi^{\nicefrac{{1}}{{3}}}\sigma^{\nicefrac{{2}}{{3}}}}{p^{\nicefrac{{1}}{{3}}}\rho^{\nicefrac{{1}}{{3}}}T^{\nicefrac{{2}}{{3}}}}+\frac{\chi}{p\rho T}\right),$
    SC: $\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{nT}+\frac{\sigma^{2}\chi}{p\rho T^{2}}+\mathrm{exp}\left[-\nicefrac{{p\rho T}}{{\chi}}\right]\right),$
    where $\sigma^{2}$ denotes the variance of the stochastic gradient, $\rho:=1-\lambda_{2}$ is a topology-dependent quantity that approaches 0 for a large and sparse network, $\chi\geq 1$. To the best of our knowledge, it is the first work that establishes the convergence rate of probabilistic distributed methods for non-convex settings.
@@
  • We elucidate the effects of noise, local steps, and data heterogeneity on the convergence of ProxSkip in stochastic non-convex, convex, and strongly convex settings. We demonstrate the robustness of ProxSkip against data heterogeneity while enhancing communication efficiency by local updates. Furthermore, we show that the convergence rates exhibited by ProxSkip in stochastic settings are comparable with those of existing state-of-the-art distributed algorithms incorporating local updates [31, 35, 36, 37, 42] (Table II).
@@ II-A Network Graph
The mixing matrix $\mathbf{W}\in[0,1]^{n\times n}$ is symmetric, doubly stochastic, and primitive. Let $\lambda_{1}=1$ denote the largest eigenvalue of the mixing matrix $\mathbf{W}$, and the remaining eigenvalues are denoted as $1>\lambda_{2}\geq\lambda_{3}\geq\cdots\geq\lambda_{n}>-1$.
@@ II-B Algorithm Description
Here, $\alpha>0$ is the stepsize, $\beta>0$, $\mathbf{G}^{t}=[\mathbf{g}_{1}^{t},\mathbf{g}_{2}^{t},\ldots,\mathbf{g}_{n}^{t}]^{\sf T}\in\mathbb{R}^{n\times d}$ with $\mathbf{g}_{i}^{t}$ representing the stochastic gradient of $\nabla f_{i}(\mathbf{x}_{i}^{t})$, $\theta_{t}=1$ with probability $p$ and $\theta_{t}=0$ with probability 1p1-p, and $\mathbf{Y}^{t}$ is the control variate. At each iteration $t\geq 0$, communication takes place with a probability $p\in(0,1]$. In the absence of communication, the update $\mathbf{X}^{t+1}=\mathbf{X}^{t}-\alpha\mathbf{G}^{t}-\mathbf{Y}^{t}$ is performed, while $\mathbf{Y}^{t+1}$ remains unchanged. This allows for multiple iterations of local computations to be performed between communication rounds. Decomposing the updates for individual nodes, we provide a detailed implementation in Algorithm 1.
@@
2: Flip coins $\theta_{0},\ldots,\theta_{T-1}]$, where $\theta_{t}\in\{0,1\}$, with $\mathop{{\bf{P}}}(\theta_{t}=1)=p$.
@@
5:  $\hat{{\bf{z}}}^{t}_{i}={\bf{x}}^{t}_{i}-\alpha{\bf{g}}^{t}_{i}-{\bf{y}}^{t}_{i}$ $\triangleright$ update the prediction variate $\hat{{\bf{z}}}^{t}_{i}$
@@
7:   ${\bf{x}}^{t+1}_{i}=\sum_{j=1}^{n}\widetilde{W}_{ij}\hat{{\bf{z}}}^{t}_{j}$ $\triangleright$ communicate with probability $p$
8:   ${\bf{y}}^{t+1}_{i}={\bf{y}}^{t}_{i}+\beta(\hat{{\bf{z}}}^{t}_{i}-{\bf{x}}^{t+1}_{i})$ $\triangleright$ update the control variate ${\bf{y}}^{t+1}_{i}$
@@
10:   ${\bf{y}}^{t+1}_{i}={\bf{y}}^{t}_{i},~{\bf{x}}^{t+1}_{i}=\hat{{\bf{z}}}^{t}_{i}$ $\triangleright$ skip communication
@@ III-A Preliminary
When $\sigma^{2}=0$, by setting $\alpha=\nicefrac{{1}}{{L}}$ and $\chi=1$, we can deduce from (3) that the communication complexity of ProxSkip to achieve $\epsilon$-accuracy, i.e., $\mathbb{E}[|\bar{{\bf{x}}}^{t}-{\bf{x}}^{\star}\|^{2}]\leq\epsilon a_{0}$, is given by $\mathcal{O}((p\kappa+\nicefrac{{1}}{{p(1-\lambda_{2})}})\mathrm{log}\nicefrac{{1}}{{\epsilon}})$, where $\kappa=\nicefrac{{L}}{{\mu}}$. If the network is sufficiently well-connected, i.e., $\nicefrac{{1}}{{(1-\lambda_{2})\kappa}}<1$, and we set $p=\sqrt{\nicefrac{{1}}{{(1-\lambda_{2})\kappa}}}$, the iteration complexity becomes $\mathcal{O}(\sqrt{\nicefrac{{\kappa}}{{1-\lambda_{2}}}}\ \mathrm{log}\nicefrac{{1}}{{\epsilon}})$, achieving the optimal communication complexity as proven by [50].
When $\sigma^{2}\neq 0$, based on (3) and the fact that $\frac{\alpha^{2}\sigma^{2}}{1-\zeta}=\frac{\alpha^{2}\sigma^{2}}{\alpha\mu}=\mathcal{O}(\alpha\sigma^{2})$, we can conclude that the local solution ${\bf{x}}_{i}^{t}$ generated by ProxSkip converges to the global minimizer $\bf{x}$$^{\star}$ at a linear rate until it reaches an $\mathcal{O}(\alpha\sigma^{2})$-neighborhood of $\bf{x}$$^{\star}$. Note that linear speedup is the property that the dominant term in the number of iterations $T$ needed to achieve $\epsilon$ accuracy is divided by $n$, e.g., $\frac{1}{n\epsilon^{2}}$ for nonconvex and $\frac{1}{n\epsilon}$ for SC, implying that increasing the number of nodes speeds up the convergence. The original result in [43] cannot choose a stepsize to achieve this due to the term $O(\alpha\sigma^{2})$ not divided by $n$. Therefore, further analysis is required to achieve the desired linear speedup.
@@ III-B Main Theorem—Convergence Rate of ProxSkip
Suppose that Assumptions 1, 2, and 4 hold. Let $\bar{{\bf{x}}}^{t}=\frac{1}{n}\sum_{i=1}^{n}{\bf{x}}_{i}^{t}$ denote the iterates of Algorithm 1 and $\bf{x}$$^{\star}$ solves (1). For sufficiently small $\alpha\leq\mathcal{O}\left(\frac{p(1-\lambda_{2})}{\chi L}\right)$ for N-C/C and $\alpha\leq\mathcal{O}\left(\frac{p\mu(1-\lambda_{2})}{\chi L^{2}}\right)$ for SC, $\chi=\mathcal{O}\left(\max\left\{1,\frac{1-p}{1-\lambda_{2}}\right\}\right)$, and $\beta=1$, we have the following convergence results.
@@
$\leq\mathcal{O}\bigg(\underbrace{\frac{F_{0}}{\alpha T}+\frac{\chi^{2}L^{2}\varsigma^{2}_{0}\alpha^{2}}{p^{2}(1-\lambda_{2})^{2}T}}_{{\mathrm{deterministic\ part}}}+\underbrace{\frac{\alpha\sigma^{2}L}{n}+\frac{\chi L^{2}\sigma^{2}\alpha^{2}}{p(1-\lambda_{2})}}_{{\mathrm{stochastic\ part}}}\bigg).$ (4)
@@
$\leq\mathcal{O}\bigg(\underbrace{\frac{R_{0}^{2}}{\alpha T}+\frac{\chi^{2}L\varsigma^{2}_{0}\alpha^{2}}{p^{2}(1-\lambda_{2})^{2}T}}_{{\mathrm{deterministic\ part}}}+\underbrace{\frac{\alpha\sigma^{2}}{n}+\frac{\chi L\sigma^{2}\alpha^{2}}{p(1-\lambda_{2})}}_{{\mathrm{stochastic\ part}}}\bigg).$ (5)
@@
$\mathbb{E}\!\left[\big\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\big\|^{2}\right]\leq$ $\underbrace{\Big(1-\frac{\alpha\mu}{4}\Big)^{T}\Big(\big\|\bar{{\bf{x}}}^{0}-{\bf{x}}^{\star}\big\|^{2}+\frac{8\chi\alpha^{2}\varsigma_{0}^{2}}{p(1-\lambda_{2})}\Big)}_{{\mathrm{deterministic\ part}}}$
$+\mathcal{O}\bigg(\underbrace{\frac{\alpha\sigma^{2}}{\mu n}+\frac{\chi L\sigma^{2}\alpha^{2}}{p\mu(1-\lambda_{2})}}_{{\mathrm{stochastic\ part}}}\bigg).$ (6)
For the non-convex setting, Theorem 1 demonstrates that the ProxSkip algorithm converges to a radius around some stationary point. Without any additional assumptions, a stationary point is the best guarantee possible and is a satisfactory criterion to measure the performance of distributed methods with nonconvex objectives [31]. For the convex case, Theorem 1 shows that ProxSkip converges around some optimal solution. When $\sigma^{2}=0$, i.e., in the deterministic case, ProxSkip converges exactly with sublinear and linear rates for N-C/C and SC settings, respectively.
Based on Theorem 1, we can even get a tighter rate by carefully selecting the stepsize to obtain the following result.
@@
Non-convex: There exists constant stepsize such that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(\bar{{\bf{x}}}^{t})\|^{2}]$
$\leq\mathcal{O}\left(\left(\frac{\sigma^{2}}{nT}\right)^{\frac{1}{2}}+\left(\frac{\chi\sigma^{2}}{p(1-\lambda_{2})T^{2}}\right)^{\frac{1}{3}}+\frac{\chi}{p(1-\lambda_{2})T}\right).$ (7)
Convex: Under the additional Assumption 3 with $\mu\geq 0$, there exists constant stepsize such that
$\leq\mathcal{O}\left(\left(\frac{\sigma^{2}}{nT}\right)^{\frac{1}{2}}+\left(\frac{\chi\sigma^{2}}{p(1-\lambda_{2})T^{2}}\right)^{\frac{1}{3}}+\frac{\chi}{p(1-\lambda_{2})T}\right).$ (8)
Strongly Convex: Under the additional Assumption 3 with $\mu>0$, there exists constant stepsize such that
$\mathbb{E}[\|\bar{{\bf{x}}}^{T}-{\bf{x}}^{\star}\|^{2}]$
$\leq\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{nT}+\frac{\sigma^{2}\chi}{p(1-\lambda_{2})T^{2}}+\mathrm{exp}\Big[-\frac{p(1-\lambda_{2})T}{\chi}\Big]\right).$ (9)
When $T$ is sufficiently large, the term $\frac{\sigma}{\sqrt{nT}}$ (and $\frac{\sigma^{2}}{nT}$ for SC) will dominate the rate. In this scenario, ProxSkip requires $T=\Omega(\frac{\sigma^{2}}{n\epsilon^{2}})$ (and $T=\Omega(\frac{\sigma^{2}}{n\epsilon})$ for SC setting) iterations to reach a desired $\epsilon$-accurate solution, thus the convergence accuracy improves linearly with $n$.
It is crucial to note that while reducing $p$ leaves the first term of the convergence rate unchanged, it cannot be reduced indiscriminately. When $p$ is above a certain threshold, the first term (independent of $p$) dominates. Reducing $p$ in this regime has minimal impact on the rate, i.e., we can skip some communication almost for free, which explains why local updates can accelerate communication. However, if $p$ falls below this threshold, the second and third terms become dominant, and any further reduction in $p$ will directly slow down convergence.
Corollary 2.
Same settings as in Theorem 1 and Corollar 1, we have the following convergence results.
@@
$p\times\underbrace{\mathcal{O}\left(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\chi}}{{\sqrt{p}}\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{\chi}{{p}(1-\lambda_{2})\epsilon}\right)}_{{\mathrm{iteration~rounds}}}$ (10)
@@
$p\times\underbrace{\mathcal{O}\left(\frac{\sigma^{2}}{n\epsilon^{2}}+\frac{\sqrt{\chi}}{{\sqrt{p}}\sqrt{1-\lambda_{2}}}\frac{\sigma}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{\chi}{{p}(1-\lambda_{2})\epsilon}\right)}_{{\mathrm{iteration~rounds}}}$ (11)
@@
$p\times\underbrace{\tilde{\mathcal{O}}\left(\frac{\sigma^{2}}{n\epsilon}+\frac{\sqrt{\chi}}{{\sqrt{p}}\sqrt{1-\lambda_{2}}}\frac{\sigma}{\sqrt{\epsilon}}+\frac{\chi\mathrm{log}\nicefrac{{1}}{{\epsilon}}}{{p}(1-\lambda_{2})}\right)}_{{\mathrm{iteration~rounds}}}$ (12)
[…]
@@
local-GT [35, 36] $\mathcal{O}\left(\frac{\sigma^{2}}{nK\epsilon^{2}}+\left(\frac{\sigma}{\rho^{2}\sqrt{K}}\right)\frac{1}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{\rho^{2}\epsilon}\right){\textsuperscript{b}}$ $\tilde{O}\left(\frac{\sigma^{2}}{nK\epsilon}+\frac{\sigma}{\sqrt{K}\rho}\frac{1}{\epsilon^{1/2}}+\frac{1}{\rho^{2}}\mathrm{log}\nicefrac{{1}}{{\epsilon}}\right)$
$K$-GT [37] $\mathcal{O}\left(\frac{\sigma^{2}}{nK\epsilon^{2}}+\left(\frac{\sigma}{\rho^{2}\sqrt{K}}\right)\frac{1}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{\rho^{2}\epsilon}\right){\textsuperscript{b}}$ no results
Periodical GT [37] $\mathcal{O}\left(\frac{\sigma^{2}}{nK\epsilon^{2}}+\left(\frac{\sigma}{\rho^{2}}\right)\frac{1}{\epsilon^{\nicefrac{{3}}{{2}}}}+\frac{1}{\rho^{2}\epsilon}\right){\textsuperscript{b}}$ no results